OpenStudy (anonymous):

6. An open box is formed by cutting squares with side lengths of 3 inches from each corner of a square piece of paper. What is a side length of the original paper if the box has a volume of 675 cubic inches? (Points : 3) 21 inches 24 inches 33 inches 36 inches

OpenStudy (anonymous):

What would be my equation?

OpenStudy (anonymous):

(x-3)*3?

OpenStudy (cwrw238):

no volume = height * area of the base

OpenStudy (anonymous):

So just solve that equation?

OpenStudy (anonymous):

So solve that equation?

OpenStudy (anonymous):

It does not make sense! it say,s x=−12,18

OpenStudy (anonymous):

Just what is the answer?

OpenStudy (anonymous):

My session is bout to end! I will give medal

OpenStudy (cwrw238):

you haven't spotted my deliberate mistake then?

OpenStudy (cwrw238):

length of side of the base = (x - 6) , height = 3 3(x -6)^2 = 675 (x - 6)^2 = 225 x-6 = 15 x = 21

OpenStudy (anonymous):

You keep deleting your answers

OpenStudy (cwrw238):

i know - that was all wrong the lengths of the base of the box = x - 3 - 3 = x - 6 - 2 lots of 3 inches have to be subtracted

OpenStudy (cwrw238):

get it?

OpenStudy (cwrw238):